A flea walking along a ruler moves from the 45 cm mark to the 27 cm mark.What is the displacement
1
Expert's answer
2019-01-24T10:23:07-0500

In accordance with the definition the displacement equals to Δx = x2 - x1 = 27 - 41 = - 14 (cm).